Recent Log Entries

MBC5S08V trim card swapped

Log Number: 4187227Logbook(s): Entry Makers: kpriceBody: 

E101B12 ch 24

fig 1: noisy/frequent mismatcher, would not cycle. Swapped with spare (fig 2) and all is well, though it should be noted that the trim rack was quite hot.

north linac gangphase drifting back and forth?

Log Number: 4187222Logbook(s): Entry Makers: kpriceBody: 

figures 1 and 2 show an adjustment from 130.8 to 130.3. It's been going back and forth around these values for a while. Figures 3 and 4 show an adjustment in the other direction (from 130.3 -> 130.6)

LERF Vault is now ODH2

Log Number: 4187201Logbook(s): Entry Makers: weaksmcBody: 

LERF Vault is now ODH2.

L2HE J23 will be brought up to 4K this afternoon, in preparation of this we are posting the LERF Vault ODH2.

Cabling of return valves on the L2HE Module pose a risk of venting reliefs, and generating an ODH hazard in the vault, any personnel accessing the vault must have appropriate ODH Training and equipment.

MYA2T01 check-up

Log Number: 4187192Logbook(s): Body: 

After getting permission and channel access from the crew chief, briefly applied .5A to MYA2T01 during a short beam outage to verify it's operation ahead of tonight's 1st pass ops. The YA responded as commanded and was then set back to zero output.

Hall C -> 30 uA on LH2 (loop 2)

Log Number: 4187172Logbook(s): Entry Makers: kpriceBody: 

Hall C requests 30 uA on LH2 (loop 2). 2x2 raster is on. NPS and correctors are off.

Beam authorization: up to 40 uA CW
Ops restrictions for this target: up to 40 uA with minimum 2x2 raster always on

Ion chambers are set for this target with 30% margin. MaxJuice is 35 uA.

iCalibrate: Hall C

Log Number: 4187158Logbook(s): Body: Calibration ParametersHall:CTarget: LH2 2 10cmPass: 5Note: Sweeper offCalibrated On:2023-Sep-17 17:22Calibrated By: dthomasFile:C-2023-09-17_1722_LH2_2_10cm_Pass-5.hcdIon ChamberExisting Setpoing (rads/hr)New Setpoint (rads/hr)Compton

COMMENT ON Impluse response testing of 2L26

Log Number: 4187150Logbook(s): Body: 

Change in plans. After talking to Rama about the scope of the task there is a concern that this could run late-ish into the evening. The new plan is to move the remote controlled hammer to the tunnel in a non-radiation area near the elevator tomorrow or in the alcove down near 2L18.

linac NDXs

Log Number: 4187148Logbook(s): Entry Makers: beneschBody: 

I rearranged gradient in NL11 yesterday to bring that trace (slate grey, upper left) down. Something turned on in 1L22 (green trace) around 1900. C100 sums, lower right of figure 2, are less than half those of the last run.

iCalibrate: Hall C

Log Number: 4187137Logbook(s): Body: Calibration ParametersHall:CTarget: Out of BeamPass: 5Note: Calibrated On:2023-Sep-15 22:55Calibrated By: anthonyFile:C-2023-09-15_2255_Out_of_Beam_Pass-5.hcdIon ChamberExisting Setpoing (rads/hr)New Setpoint (rads/hr)Compton3,000<

BEM BPM Status/Connectivity intermittent

Log Number: 4187129Logbook(s): Entry Makers: lrichardBody: 

This has been happening for the East Arc BPMs. Master Switch cycle reconnects them but they should not lose connection as often as they have been lately. Contacted EESIC and they will look into it. Wanted to put something in the log about it. Spoke with Software Group and they said BEM gets the BPM signal from the High Level Status bit (HLST). That might be the first thing to investigate...

Pages